# Detection of pairwise kSZ effect with DESI galaxy groups and Planck in Fourier space

### 07 Jan 2024 v1

@@ Appendix C Masked kSZ dipole with higher order expansion
We compare the evaluated $\tilde{P}_{\rm kSZ}^{\ell=1}$ from the above equation with that from eq. (43), and the result is shown on the left panel of figure 13. As illustrated, the difference between eq. (43) and eq. (C1) is less than $0.1\%$ and in turn, it makes sense to ignore the higher multipoles of the kSZ correlation function in evaluating the theoretical masked kSZ dipole.
@@ Appendix D Results of group samples with different masses
In this work we select the group sample by its lower mass threshold $M_{\rm th}$, based on the fact that $\bar{\tau}\propto M$, as validated by figure 8. By varying $M_{\rm th}$ and $\theta_{\rm AP}$, the measurement S/Ns of different samples with various AP filter radii are shown on the right panel of figure 14. We choose the sample with the highest S/N as the baseline sample, which includes $5.65\times 10^{6}$ heaviest groups. On the left panel of figure 14, the $\bar{\tau}$ profiles of different group samples are shown, by which the mass dependence of the gas distribution within and around dark matter halos will be studied in a companion paper.

### 01 Mar 2024 v2

@@ Appendix C Masked kSZ dipole with higher order expansion
We compare the evaluated $\tilde{P}_{\rm kSZ}^{\ell=1}$ from the above equation with that from Equation (43), and the result is shown on the left panel of figure 13. As illustrated, the difference between Equation (43) and Equation (C1) is less than $0.1\%$, and in turn, it makes sense to ignore the higher multipoles of the kSZ correlation function in evaluating the theoretical masked kSZ dipole.
@@ Appendix D Results of group samples with different masses
In this work, we select the group sample by its lower mass threshold $M_{\rm th}$, based on the fact that $\bar{\tau}\propto M$, as validated by figure 8. By varying $M_{\rm th}$ and $\theta_{\rm AP}$, the measurement S/Ns of different samples with various AP filter radii are shown on the right panel of figure 14. We choose the sample with the highest S/N as the baseline sample, which includes $5.65\times 10^{6}$ heaviest groups. On the left panel of figure 14, the $\bar{\tau}$ profiles of different group samples are shown, by which the mass dependence of the gas distribution within and around dark matter halos will be studied in a companion paper.